How would you write this rational expression in lowest terms?
6-x/x-6

Found 2 solutions by checkley71, marcsam823:
Answer by checkley71(8403)

IF YOU MEAN 6-(X/X)-6 THEN 6-1-6=-1
IF YOU MEAN (6-X)/(X-6)THEN THIS ALSO =-1
PROOF:
SET X=ANY NUMBER & SOLVE.
EXAMPLE X=9.
(6-9)/9-6)=-3/3=-1

Answer by marcsam823(57)

1. Factor out -1 from the numerator
-1(x-6)/(x-6)
2. Cancelling out the (x-6) terms leaves you with the lowest terms result of:
-1